Story problem help.
ellis hits a baseball the path of the ball is approximated by: y = - 1/500 (x-200)^2 +76
I have solved some of the questions - baseball path vertex (200, 76)
Find the height of the ball after it traveled 220 feet horizontally.
-rewrite the equation above in standard form.

Answer by lwsshak3(11628)

ellis hits a baseball the path of the ball is approximated by: y = - 1/500 (x-200)^2 +76
I have solved some of the questions - baseball path vertex (200, 76)
Find the height of the ball after it traveled 220 feet horizontally.
-rewrite the equation above in standard form.
y = - 1/500 (x-200)^2 +76
f(200)=-1/500(220-200)^2+76
f(200)=-1/500(20)^2+76
f(200)=-1/500(400)+76=-4/5+76=75.2
height of the ball after it traveled 220 feet horizontally=75.2 ft
...
standard(general) form of equation:y=Ax^2+Bx+c
y = - 1/500 (x-200)^2 +76
500y=-(x^2-400x+40000)+38000
500y=-x^2+400x-40000)+38000
500y=-x^2+400x-2000
y=-(1/500)x^2+0.8x-4
...
y-intercept
set x=0
y=-4
..
x-intercepts:
x≈5.0641 & 394.396
..
ball will hit the 15 ft fence about 8 ft from the bottom